OpenStudy (anonymous):

find an equation of the tangent line to the hyperbola y=3/x at the point (3,1) Thanks

OpenStudy (dumbcow):

dy/dx = slope of tangent line line goes through point (3,1) \[\frac{d}{dx} \frac{3}{x} = \frac{d}{dx}3x^{-1} = -3x^{-2} = -\frac{3}{x^{2}}\] x=3 --> slope = -1/3 \[y-1 = -\frac{1}{3}(x-3)\] \[y = -\frac{1}{3}x +2\]

OpenStudy (anonymous):

thank you :)